How many grams of chocolate worth P600 per kg must be mixed with 10 jg of chocolate worth P450 per kilogram to produce a moxture worth P500 per kilogram?
Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(52781)

.


a TWIN problem is just solved 10 minutes ago under this link

https://www.algebra.com/algebra/homework/word/mixtures/Mixture_Word_Problems.faq.question.1173323.html

https://www.algebra.com/algebra/homework/word/mixtures/Mixture_Word_Problems.faq.question.1173323.html


It is your TEMPLATE.

Having the template in front of you, solve the current problem in the same way.



Answer by greenestamps(13200)



Here is a quick and easy way to solve 2-part "mixture" problems like this without formal algebra (if, of course, a formal algebraic solution is not required).

The price of the mixture per kilogram, P500, is 1/3 of the way from P450 to P600, the prices per kg of the two ingredients. (look at it as 500 begin 1/3 of the way from 450 to 600 on a number line....)

That means 1/3 of the mixture is the more expensive ingredient.

That means the 10kg of the less expensive chocolate is 2/3 of the mixture; and that mean 1/3 of the mixture is 5kg.

ANSWER: 5kg

CHECK:
10(450)+5(600) = 4500+3000 = 7500
15(500) = 7500